OpenStudy (anonymous):

The tables below show the values of f(x) and g(x) for different values of x. f(x) = 2(3)x x f(x) -2 0.22 -1 0.67 0 2 1 6 2 18 g(x) = 3x + 9 x g(x) -2 9.11 -1 9.33 0 10 1 12 2 18 Based on the tables, what is the solution to the equation 2(3)x = 3x + 9? x = 0 x = 2 x = 12 x = 18

OpenStudy (anonymous):

@halorazer any idea bro?

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

@iambatman @Loser66 @Compassionate

OpenStudy (anonymous):

The tables don't make any sense. in g(x)=3x+9, if we took the values given and put them in, they are not equal. \[9.11=3 \times -2 +9\] \[9.11=-6+9\] \[9.11=3\] Which is simply not true.

OpenStudy (anonymous):

if u had 2 guess wat wuld u put ;)

OpenStudy (anonymous):

Sir, if you could not wink at me, that would be fabulous.

OpenStudy (anonymous):

But, given what is in the tables, because when x is 2, both of them have a result of 18, I'd say that your answer is B.

OpenStudy (anonymous):

:I

OpenStudy (anonymous):

ty xD

OpenStudy (anonymous):

wana help with 1 more o_O

OpenStudy (anonymous):

Will there be winking? >.>

OpenStudy (anonymous):

ofc not :)

OpenStudy (anonymous):

Better not be. Ask away.

OpenStudy (anonymous):

The graph below shows two functions. function f of x is a straight line which joins the ordered pairs negative 3.5, 15 and 4, 0. Function g of x is a curved line which joins the ordered pairs negative 1.7, 11.5 and 4, 0 Based on the graph, what are the approximate solutions to the equation –2x + 8 = (0.25)x? 8 and 4 1 and 4 –1.7 and 4 1.7 and –4

OpenStudy (anonymous):

@halorazer any idea brahhh?

OpenStudy (anonymous):

I have every idea. But this one is way too simple for me to just give you the answer.

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

The solutions they are referring to are where the two lines intersect, which is in two different places. Those are your answers.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/5366cb14e4b01162a64f5afb cum on here

OpenStudy (anonymous):

ill give u another medal o_O

OpenStudy (anonymous):

if u can help :O

OpenStudy (anonymous):

yay